Solve for x
-5by/6(y+w-z)=x-6ax

Answer by stanbon(75887)

Solve for x
-5by/6(y+w-z)=x-6ax
-5by/6(y+w-z)=x(1-6a)
x = -5by/[6(1-6a)(y+w-z)]
=============================
Cheers,
Stan H.